### v2

# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

Abstract.
Recommender systems play a pivotal role in helping users navigate an overwhelming selection of products and services. On online platforms, users have the opportunity to share feedback in various modes, including numerical ratings, textual reviews, and likes/dislikes. Traditional recommendation systems rely on users’ explicit ratings or implicit interactions (e.g., likes, clicks, shares, saves) to learn user preferences and item characteristics. Beyond these numerical ratings, textual reviews provide insights into users’ fine-grained preferences and item features. Analyzing these reviews is crucial for enhancing the performance and interpretability of personalized recommendation results. In recent years, review-based recommender systems have emerged as a significant sub-field in this domain. In this paper, we provide a comprehensive overview of the developments in review-based recommender systems over recent years, highlighting the importance of reviews in recommender systems, as well as the challenges associated with extracting features from reviews and integrating them into ratings. Specifically, we present a categorization of these systems and summarize the state-of-the-art methods, analyzing their unique features, effectiveness, and limitations. Finally, we propose potential directions for future research, including the integration of multimodal data, multi-criteria rating information, and ethical considerations.
Review-based recommender systems, Deep Learning, Contrastive Learning, Multimodal Learning, LLMs
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: XXjournalnumber: Xarticle: XXXpublicationmonth: 4ccs: Recommender Systemccs: Customer review
1. Introduction
In the rapidly advancing digital era, the sheer volume of online content, including products, movies, restaurants, and services, presents a significant challenge: information overload. This overwhelming amount of information often makes it difficult for people to find what they truly need or prefer. In response, recommender systems or information filtering techniques have emerged as key solutions, tackling the issue of consumer choice by guiding users to content that aligns with their interests and preferences. These systems play an important role in providing personalized recommendations, enhancing user experiences, contributing to business growth, and simplifying the decision-making process. These systems are extensively implemented by major businesses such as Google, Facebook, and Amazon, and across a wide range of web domains, including online shopping platforms (Schafer et al., 1999; Hwangbo et al., 2018; Zhu et al., 2014), healthcare services (Tran et al., 2021), educational resources (Tan et al., 2008), news portals (Raza and Ding, 2019a), and media streaming services (Majid et al., 2013).
Collaborative Filtering (CF) (Su and Khoshgoftaar, 2009) and Content-based Filtering (CBF) (Thorat et al., 2015) have been fundamental to developing personalized recommender systems. CF excels when ample user ratings are available but struggles with sparse data, leading to inaccuracies and a lack of transparency. Content-based approaches aim to mitigate some of CF’s limitations by representing users and items through various data forms, focusing on item attributes and user preferences. Despite these efforts, content-based approaches encounter their own challenges due to the dynamic nature of user interests and content. With the proliferation of review data in online platforms, a new avenue for recommendation systems has emerged. Unlike traditional recommender systems that primarily rely on numerical ratings or browsing history, review-based systems explore the rich textual feedback provided by users. Review-based systems (Wu et al., 2019a; Liu et al., 2019a; Zheng et al., 2017; Chen et al., 2015; Seo et al., 2017; Chin et al., 2018; Huang et al., 2020a; Al-Ghuribi and Noah, 2019) leverage the wealth of information contained within user reviews, offering a promising solution to the limitations of the traditional rating-based methods.
Review-based recommender systems harness the power of user reviews, extracting fine-grained aspects and associated sentiments, to offer more personalized, accurate, and explainable recommendations (Wu et al., 2019a). However, incorporating user reviews into recommendation models brings unique challenges. These include extracting fine-grained user preferences, and item characteristics from reviews, and leveraging them to learn the overall preferences. In addition, handling the vast and varied nature of textual data, and addressing issues like bias and privacy concerns are important (Raza et al., 2024). Hence, review-based recommender systems require advanced methodologies and innovative solutions.
Over the past years, extensive research has been conducted on recommender systems, including various systematic reviews (Adomavicius and Tuzhilin, 2005; Konstan and Riedl, 2012; Burke, 2002; Shi et al., 2014; Jannach et al., 2015; Khan et al., 2017; Batmaz et al., 2019). However, there appears to be a gap in comprehensive reviews specifically targeting review-based recommender systems. Previous surveys have broadly covered traditional recommender systems, deep learning based systems (Zhang et al., 2019), recommender systems using graph neural networks (Gao et al., 2023; Wu et al., 2022), multi-criteria recommender systems (Al-Ghuribi and Noah, 2019), session-based recommender systems (Wang et al., 2021a) and more recently reinforcement learning-based recommendation systems (Afsar et al., 2022), but have not concentrated exclusively on review-based systems. Bobadilla et al. (Bobadilla et al., 2013) presents an extensive overview of recommender systems and collaborative filtering, discussing their evolution and future directions. Lu et al. (Lu et al., 2015) surveyed application developments in recommender systems, categorizing applications and summarizing key algorithms. There are also survey papers that focus on context-aware recommendation systems, offering overviews of recent developments and future challenges (Haruna et al., 2017; Raza and Ding, 2019b).
A significant insight on deep learning-based recommender systems are presented in (Zhang et al., 2019), summarizing the state-of-the-art developments and current trends in the field. The study highlights the advantages and disadvantages of deep learning applications in this area. A comprehensive survey on deep learning focused knowledge graph is presented in (Gao et al., 2020a). The paper exclusively emphasizes on state-of-the-art frameworks (e.g. graph embeddings), and addresses the challenges such scalability, and cold-start issues (Gao et al., 2020a). This survey also summarizes commonly used datasets, evaluation methods, and open-source codes, while proposing potential research directions. More recently, the latest developments and trends in recommendation systems have been explored to include the consideration of robustness, data bias, and fairness, alongside with a brief review of personalized and group recommender systems (Li et al., 2023a)
Despite these extensive surveys, the specific area of review-based recommender systems is overlooked. This paper aims to fill this gap by focusing exclusively on review-based recommender systems. Here, we explore various methods for feature extraction from reviews, categorize advanced techniques, identify key challenges and solutions specific to this domain. We also discuss future possibilities in review-based recommender systems, offering insights into their development, methodologies, and potential advancements. Thus, our survey, with its focused approach, stands out from previous general surveys by providing in-depth insights into the rapidly evolving field of review-based recommender systems.
Contribution of this survey: The key contributions of this survey can be summarized as follows:
  1. (1)
    We systematically explore review-based recommender systems and introduce a classification scheme that considers feature extraction techniques and methodologies used to incorporate reviews into the recommendation process. To the best of our knowledge, this is the first comprehensive survey in this particular domain.
  2. (2)
    We provide a detailed overview of the state-of-the-art models, analyzing their unique features, effectiveness, limitations, and evaluation methods.
  3. (3)
    We investigate significant challenges and recent developments, proposing innovative directions for future research that address practical and theoretical gaps in the field. Our findings suggest pathways for developing more robust and transparent systems.
Methodology: To conduct a systematic review on review-based recommender systems, we collect the relevant papers from a wide range of venues including ACM conference proceedings (e.g., RecSys, SIGIR, CIKM, AAAI, WWW, WSDM, KDD, UMAP, EMNLP, ACL), Google Scholar, and IEEE Xplore. Our review period spans from January 2015 to January 2024. We use specific search keywords that include ”review-based recommender systems”, ”review-aware recommender systems”, ”aspect-based recommender systems”, and ”aspect-aware recommender systems”. For conference proceedings, we initially select all papers that contain any of the above mentioned keywords in their titles, resulting in a total of 14,862 papers obtained across all conferences. In contrast, for Google Scholar and IEEE Xplore, we Utilize different strategy e.g. Boolean Operators. Our search queries are structured as (”review-based” OR ”review-aware” OR ”aspect-based” OR ”aspect-aware”) AND ”recommender systems”. Initial search results reveal a total of 5140 papers.
In contrast, IEEE Xplore reports much lower counts for similar keywords, e.g 855 papers. We perform a preliminary screening based on titles and abstracts, focusing on relevance to our study’s goals and careful selection. This screening involves two reviewers to enhance the reliability of the selection process, with discrepancies resolved through discussion, ultimately shortlisting 71 papers for detailed analysis.
These selected papers undergo a thorough content analysis to extract information on methodologies, findings, and trends’ evolution over time, using both qualitative and quantitative techniques to ensure a robust interpretation of the literature. Figure 1 illustrates the distribution of the selected publications across different sources and formats, as well as their annual distribution, helping to understand the focus and growth of research areas within the field. Considering the rapid advancements in review-based recommender systems, we plan to update our literature review annually on arXiv, incorporating recent publications to stay updated with the latest developments.
Figure 1. Trends in Review-Based Recommender Systems Research from 2015 to 2024
2. Background
Before further exploring the review-based recommender systems, we will discuss each of these categories in detail. Recommender systems are broadly classified into three categories: Collaborative Filtering (Schafer et al., 2007), Content-based Filtering (Pazzani and Billsus, 2007), and hybrid systems (Thorat et al., 2015). In the following subsections, we briefly discuss each of the categories.
2.1. Collaborative Filtering
Collaborative Filtering (CF) (Schafer et al., 2007) is a method used in recommender systems to predict a user’s interests based on the preferences of a larger user group. This technique relies on collecting and analyzing data on user behaviors, activities, or preferences to make personalized recommendations (Su and Khoshgoftaar, 2009). The core idea behind collaborative filtering is that, if two users have demonstrated similar preferences in their interactions with items in the past, it is likely that their future interaction patterns will also be similar. CF establishes a relationship between two different kinds of entities: users and items.
CF can be categorized into two main types: Memory-based (Deshpande and Karypis, 2004; Pu et al., 2011), and Model-based approaches (Paterek, 2007; Sarwar et al., 2001). Memory-based approaches make predictions about a user’s interests based on the similarity between users or items. This involves directly using the raw data of past user interactions to predict future preferences. On the other hand, model-based approaches involve building and training a model using user-item interaction data. These models, once trained, can predict a user’s interest in an item based on the learned patterns. The prediction in CF systems is primarily based on the historical data of users and their interactions with items. In memory-based methods, predictions are made by averaging the ratings of similar items or by considering the ratings of similar users. Model-based approaches, however, predict user preferences by learning the latent factors of users and items from the interaction data, often employing techniques like matrix factorization (Kalman, 1996).
Recently, the trend in CF has shifted towards more sophisticated model-based approaches, incorporating advanced machine learning and deep learning models. Neural Collaborative Filtering (NCF) models (Chen et al., 2018a), for instance, use neural networks to learn the complex non-linear relationships between users and items.
2.2. Content-based Filtering
Content-based filtering (CBF) (Pazzani and Billsus, 2007) recommends items by analyzing their attributes and comparing them to a user’s preferences or past interactions. This technique creates user profiles based on demographic information or previous interactions and aligns these with detailed item descriptions to personalize recommendations. It addresses the cold start problem effectively by suggesting items based on their features, even without prior user interaction history. While adept at recommending items similar to those a user has previously shown interest in, CBF can sometimes lack diversity in its suggestions. To overcome this, recent advancements leverage machine learning and natural language processing to refine user-profiles and item descriptions, aiming to improve the relevancy and accuracy of recommendations. Despite its challenges, CBF remains a key method for offering personalized recommendations, with continuous enhancements driven by technological progress.
2.3. Hybrid Recommender Systems
Hybrid recommender systems combine elements from both collaborative filtering and content-based filtering to offer more accurate and diverse recommendations. By integrating these approaches, hybrid systems can leverage the strengths and mitigate the weaknesses of each method. They use a variety of techniques, including combining separate collaborative and content-based models, incorporating features of one method into the other, or building a unified model that simultaneously learns user and item representations from both user interactions and item attributes. This approach allows hybrid systems to provide personalized recommendations like content-based methods while also benefiting from the broad recommendation capabilities of collaborative filtering. Hybrid systems are particularly effective in addressing the cold start problem for new users and items, as they can recommend items based on content when user interaction data is sparse. However, the hybrid recommender system is complex and expensive to implement (Su and Khoshgoftaar, 2009).
3. Why adding review into recommender systems
The ultimate goal of a recommender system is to suggest items that match user’s interests. Traditional recommender systems learn the user’s preferences based on historical rating data. However, these rating-based systems encounter significant challenges, including data sparsity, a lack of explainability, and reduced accuracy. Incorporating additional forms of information, such as text and images, is advantageous when available. Reviews, in particular, play a crucial role in understanding user preferences and item characteristics. User reviews, being free-form text, allow users to express their opinions freely in any direction. Utilizing reviews helps address these critical issues in recommender systems effectively.
Addressing Data Sparsity: Data sparsity poses a significant challenge in recommender systems, particularly in real-world scenarios where rating data is scarce, often accompanied by few reviews. Learning user preferences and item features from such sparse data is challenging. Reviews play a crucial role in addressing this challenge. Through textual feedback, users offer detailed opinions, making it valuable for extracting features to understand user preferences. Various existing recommender systems demonstrate the effectiveness of leveraging reviews in mitigating data sparsity issues (Zheng et al., 2017; Tan et al., 2016).
Improving Accuracy: Improving accuracy is another challenging task for any recommender system. Recommendation accuracy largely depends on learning user and item representation. Since reviews contain fine-grained opinions, utilizing them for feature learning can be beneficial. There has been a significant amount of work available that utilizes review to improve accuracy (Kim et al., 2016; Zhang et al., 2017; Tan et al., 2008; Seo et al., 2017). Different NLP techniques are used to extract the features from user reviews and item reviews, followed by an interaction technique to predict the ratings. Some paper utilizes both review and rating for better learning the user and item features as well as improving recommender accuracy (McAuley and Leskovec, 2013).
Enhancing Explainability: Explainability is a significant issue in recommender systems. A recommender system needs to be able to provide a rationale for its recommendations to a user. Traditional rating-based recommender system lacks interpretability because they solely rely on latent factor learning. Latent factors represent the hidden aspects that influence user preferences, but they usually cannot explain why a user gave a specific rating to an item. Reviews can be analyzed to identify the aspects or features that motivate the users to provide specific ratings. Many existing models have demonstrated that utilizing reviews can improve explainability (McAuley and Leskovec, 2013; Seo et al., 2017).
4. Review-based recommender systems
The three main components of recommender systems are user/item representation learning, user-item feature interaction mechanism, and overall rating or preference prediction. Traditional recommender systems take user $ID$ and item $ID$ as input and learn their features using latent factor models (Koren et al., 2009). However, these static feature learning methods suffer from data sparsity and scalability problems. It is often beneficial to use as much as information possible including the user review text(Chin et al., 2018), images (Yilma and Leiva, 2023), numerical rating score(Cheng et al., 2018a), video (Mu and Wu, 2023), and audio (Bogdanov et al., 2013). Many e-commerce websites such as Amazon.com  [^1] https://www.amazon.com/, Yelp.com [^2] https://www.yelp.com/, and TripAdvisor.com [^3] https://www.tripadvisor.com/ collect user feedback as textual reviews in addition to numerical ratings. User reviews are considered as a rich source of feedback for learning user preferences or item attributes. Because users are free to express their opinions, it is assumed that the users’ fine-grained preferences are available in the review. These reviews are utilized in various ways to solve problems in traditional rating-based recommender systems, e.g. data sparsity problems, and explainability problems.
Figure 2. Sample Hotel Review Highlighting Key Aspects
Figure 3. The detailed workflow of the review-based recommender systems. The top panel represents the Review-Only Recommender Systems, and the bottom panel represents the Rating and Review-Integrated Recommender Systems
We categorize the review-based recommender systems into two categories: 1) General review-based method, and 2) Aspect-based method. General review-based methods (Musat et al., 2013; Liu et al., 2019a; Zheng et al., 2017) learn overall user preferences and item characteristics from the review without considering user sentiment on any particular topic or aspects of the items. In contrast, aspect-based methods (Chin et al., 2018) rely on estimating fine-grained sentiment on individual aspects of items. The details for each of the categories are described in the subsequent subsections.
4.1. General Review-based Recommender Systems
To illustrate this category, suppose a user set $U=\{u_{1},u_{2},...,u_{|U|}\}$, item set $I=\{i_{1},i_{2},...,i_{|I|}\}$, review set $E=\{e_{1},e_{2},...,e_{|E|}\}$, and an overall rating set $R=\{r_{1},r_{2},...,r_{|R|}\}$.
In a traditional recommender system, user $u^{\prime}s$ preference $\hat{r}$ towards item $i$ is estimated as
(1) $\hat{r}=P_{u}^{\mathsf{T}}Q_{i}+b$
where $P_{u}$ and $Q_{i}$ are the latent feature vector of user $u$ and item $i$ learned using latent factor models, e.g. matrix factorization, and $b$ is the bias term.
The review-based system focuses on integrating review features with rating features to estimate the overall rating. Two different techniques are used to learn user and item representation from reviews: 1) learning user and item representation from review only, and 2) utilizing both reviews and ratings. The former focuses on creating user and item documents to learn user and item representation while the latter emphasizes integrating features from both reviews and ratings. Let’s consider the user and item representation from reviews are $X_{u}$ and $X_{i}$ respectively. Then the equation (1) for the above-mentioned two methods can be rewritten as
(2) $\hat{r}=X_{u}^{\mathsf{T}}X_{i}+b$
and
(3) $\hat{r}=concat(P_{u},X_{u})^{\mathsf{T}}concat(Q_{i},X_{i})+b$
Although the general review-based recommender system demonstrates performance improvement over the rating-only recommender system, it often lacks explainability. This is because the general recommender system focuses on learning the overall preferences for items. To make the system more trustable, it is important to consider the detailed analysis of the review and generate the explanation when recommending particular items.
4.2. Aspect-based Recommender Systems
The aspect-based recommender system aims to understand user’s fine-grained opinions on items. It is done by extracting or learning latent representations of aspects from the reviews. Consider an aspect set $A=\{a_{1},a_{2},...,a_{|A|}\}$ where $|A|$ is the number of aspects and the aspect-based representation of user and item are $X_{U}^{A}$ and $X_{I}^{A}$ respectively. The equation (2) and (3) can be reformulated as
(4) $\hat{r}=(X_{u}^{A})^{\mathsf{T}}X_{i}^{A}+b$
and
(5) $\hat{r}=concat(P_{u},X_{u}^{A})^{\mathsf{T}}concat(Q_{i},X_{i}^{A})+b$
where $X_{u}^{A}$ and $X_{i}^{A}$ are aspect level representation of users and items respectively. The publications of different categories of review-based papers are listed in Table 1.
The term “aspect” refers to the specific attributes of products. Naturally, a user might not be interested in all aspects of a product but one or more of them. Furthermore, the significance of an aspect to a user can differ from one item to another. Hence, it is vital to capture the aspects and their relevance to users. Aspect-aware recommender models leverage advanced NLP techniques such as topic modeling, deep neural networks, and attention mechanisms to dynamically model user preferences. In Figure 2 “Location”, “service”, and “cost” represent the aspects.
Table 1. Publications of different Categories of review-based papers
Categories Publications
General Review-based (Wei et al., 2024; Chiang et al., 2023; Yang et al., 2023; Xv et al., 2022; Choi et al., 2022; Shuai et al., 2022; Luo et al., 2021; Tan et al., 2021; Xiong et al., 2021; Wang et al., 2021b; Liu et al., 2021d; Gao et al., 2020b; Peña et al., 2020; Dong et al., 2020; Ni et al., 2019; Xia et al., 2019; Wu et al., 2019c; Liu et al., 2019a, b; Rafailidis and Crestani, 2019; Alexandridis et al., 2019; Deng et al., 2018; Wang et al., 2018; Hyun et al., 2018; Chen et al., 2018a; Lu et al., 2018; Zhang et al., 2017; Seo et al., 2017; Catherine and Cohen, 2017; Zheng et al., 2017; He et al., 2015; Almahairi et al., 2015; Liu et al., 2022a; Zhang et al., 2021; Lyu et al., 2021; Wu et al., 2019b; Chen et al., 2019; Mei et al., 2018; Tan et al., 2016; Kim et al., 2016; Wei et al., 2023; Xi et al., 2021; Sayyadpour and Nazarizadeh, 2023; You et al., 2023; Ferdiana et al., 2022; Huang, 2022; Cai et al., 2022; Wu et al., 2021; Du et al., 2021; Liu et al., 2021a, b, c; Shamir et al., 2021; Ray et al., 2020; Huang et al., 2020b; Peng et al., 2020; Khosla et al., 2020; Jin et al., 2020; Suzuki et al., 2019; Cao et al., 2019; Shalom et al., 2019; Li et al., 2019; He et al., 2018; Wu et al., 2019a; Zhang et al., 2016)
Aspect-based (Wang et al., 2023b, a; Yang et al., 2021; Guan et al., 2019; Cheng et al., 2018b, a; Bauman et al., 2017; Li et al., 2015; Shi et al., 2022; Chin et al., 2018; Sun et al., 2021; You et al., 2020; Da’u et al., 2020; Mowlaei and Keyvanpour, 2020; Tewari et al., 2019; Sangeetha et al., 2017; Qiu et al., 2016; Bhojwani et al., 2023; Won et al., 2022; Ozcan et al., 2022; Peng et al., 2022; Cantador et al., 2022; Liu et al., 2022b; V and Shah, 2022; Godakandage et al., 2020; Da’u and Salim, 2019; Wu and Ester, 2015; Safar et al., 2022)
Each of these categories can further be subdivided into two classes: a) Review-Only Recommender Systems, and b) Rating and Review-Integrated Recommender Systems.
Review-Only Recommender Systems: Within this classification, recommender systems utilize reviews exclusively to understand user behavior and item attributes. In this case, user and item reviews are used to construct user and item profiles, resulting in improved representation of user-item interactions. This is illustrated in the upper panel of Figure 3.
Rating and Review-Integrated Recommender Systems: Both the rating and review-based features are bridges to obtain a rich representation of the user and item. The combined strength of rating-based and review-based features is harnessed to improve feature representation. This category is depicted in the lower panel of Figure 3.
5. State-of-the-art Review-aware Recommender Systems
In this section, we briefly present the state-of-the-art review-based recommender systems. Our analysis emphasizes on the methodology, the rationale behind the method, features, datasets, evaluation metrics, and performance. We classify these systems based on the methods used into two main categories: 1) Probabilistic Approach, and 2) Deep Learning-based Approach.
5.1. Probabilistic Approaches
The main focus of the probabilistic approaches is to leverage the topic or aspect features learned from reviews and integrate them with the rating-based features, enhancing the overall recommendation accuracy, scalability, and interpretability. Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and Gaussian Mixture Models (GMM) (Reynolds et al., 2009) are the two most successful techniques utilized to achieve this goal.
Most probabilistic approaches, including the Hidden Factors as Topics (HFT) (McAuley and Leskovec, 2013) and Rating-Boosted Latent Topics (RBLT) (Tan et al., 2016), share a common methodology that integrates the latent dimension in ratings with topics extracted from reviews. This hybrid approach is particularly effective in addressing data sparsity by enhancing the system’s ability to make accurate predictions even with limited user history. Both models leverage textual reviews to refine the recommendation process, achieving improved effectiveness across diverse datasets such as Amazon and Yelp.
One of the unique features of HFT (McAuley and Leskovec, 2013) is its ability of extracting the topics from the review that can explain the variation present in ratings and reviews. In addition, this model demonstrates rating prediction accuracy, particularly for new products and users with limited rating history, compared to matrix factorization and LDA-based models. It can also identify the most useful and informative reviews, ensuring interpretability. However, it faces scalability challenges due to its computational demands. In contrast, RBLT (Tan et al., 2016) bridges the advantage of the latent factor model and topic model to integrate the ratings and textual reviews, enhancing the recommendation accuracy. Similar to HFT, RBLT alleviates data sparsity and improves interpretability by exploring item recommendability and user preference distributions in a shared topic space.
The Neural Gaussian Mixture Model (NGMM) (Deng et al., 2018) uses a Gaussian mixture model in its architecture to refine rating predictions. This model utilizes neural networks to represent user and item characteristics through latent vectors, capturing the details of preferences and properties. Its use of a Gaussian Mixture Model at a shared layer enables precise rating predictions, enhancing the integration of textual review analysis into recommender systems for improved accuracy.
Similarly, the Aspect-Aware Latent Factor Model (ALFM) (Cheng et al., 2018b) correlates latent topics extracted from reviews with latent factors from ratings, focusing on identical aspects. It surpasses traditional models by avoiding direct one-to-one correspondence between topics and factors, which improves both accuracy and interpretability but depends heavily on the quality of the review content. The Matching Distribution by Reviews (MDR) (Shalom et al., 2019) model effectively combines deep learning with collaborative filtering for text processing, capturing the distribution of user-item matches found in reviews. It provides better prediction accuracy and reduced runtime across multiple datasets. Finally, the Topic Initialized Latent Factor Model (TIM) (Peña et al., 2020) enhances recommendations accuracy by initializing user and item embeddings within a topic space derived from reviews. This method not only speeds up model convergence but also improves interpretability by aligning embeddings with identifiable topics, offering a more transparent and understandable recommender system.
While all these models demonstrate enhanced recommendation accuracy, their effectiveness varies by application domain. For instance, HFT and RBLT are particularly beneficial for new products and users in e-commerce, where user history might be limited. In contrast, NGMM and the ALFM show broader applicability by also enhancing recommendations in specialized markets and for unique item groups. The detailed summary of the probabilistic models is presented in Table 2.
Figure 4. Summary of the different review-based recommender models
Table 2. The detailed summary of the Probabilistic Approaches
Category Models Year Architecture Features Datasets Evaluation Tasks
Probabilistic HFT(McAuley and Leskovec, 2013) 2013 LDA ratings & reviews amazon, yelp & beer MSE & MAE accuracy & interpretability
RBLT (Tan et al., 2016) 2016 LDA ratings & reviews amazon MSE, P & R accuracy & interpretability
ALFM(Cheng et al., 2018b) 2018 topic reviews amazon & yelp RMSE accuracy
NGMM(Deng et al., 2018) 2018 NN, GMM reviews amazon MSE feature extraction
MDR (Shalom et al., 2019) 2019 LDA reviews amazon & yelp MSE accuracy & interpretability
TIM (Peña et al., 2020) 2020 LDA, MF ratings & reviews amazon & tripAdvisor Recall accuracy & interpretability
5.2. Deep Learning-based Approaches
Deep learning, a subset of machine learning, involves training large neural networks to recognize patterns in data. Deep learning architectures such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, GRU (Gated Recurrent Unit), Attention Mechanism, GNN (Graph Neural Networks), and Autoencoders have revolutionized various fields, including image and speech recognition, natural language processing, and recently, recommender systems. In recommender systems, deep learning-based models have shown an exceptional ability to extract deep, non-linear relationships between users and items, significantly improving recommendation quality. These models can handle vast amounts of data, learning from user interactions, item attributes, and additional context to provide personalized recommendations.
Review-based recommender systems, in particular, have benefited from deep learning models. By analyzing the rich textual data in reviews, these systems uncover nuanced user preferences and item features that traditional methods might miss. In this subsection, we briefly summarize state-of-the-art deep learning-based models and analyze the key features along with their effectiveness. Some papers may apply multiple deep learning techniques in their recommender systems, and in this case, we categorize them based on the main models they have used in their systems.
Convolutional Neural Network-based Approaches: Convolutional Neural Network (CNN) based models typically excel in addressing challenges associated with sparse user-to-item interactions by extracting the contextual features from textual reviews. CNNs can identify and capture local features, such as phrases and sentences, that carrys significant semantic meaning. This capability allows these models to make more accurate predictions, even with limited data. Most CNN-based approaches, such as Convolutional Matrix Factorization (ConvMF) (Kim et al., 2016) and Deep Cooperative Neural Networks (DeepCoNN) (Zheng et al., 2017), combine traditional Matrix Factorization methods with CNN’s power to enhance the overall effectiveness of recommender systems.
The Convolutional Matrix Factorization (ConvMF) (Kim et al., 2016) integrates CNNs with Probabilistic Matrix Factorization to effectively capture detailed context within item descriptions, moving beyond the limitation of traditional bag-of-word techniques. ConvMF is particularly noted for its ability to improve recommendation accuracy in conditions of extreme data sparsity. However, its complexity and the detailed benefits of using pre-trained word embeddings indicate potential areas for further refinement and optimization. Utilizing dual neural networks, the Deep Cooperative Neural Networks (DeepCoNN) (Zheng et al., 2017) learns item properties and user behaviors from review texts. Its shared layer mimics factorization machine techniques, allowing intricate interactions of latent factors between users and items. Expanding on the concept of network representation, the Translation-Based Network Representation Learning (TransNet) (Catherine and Cohen, 2017) introduces vertex interactions in network representation learning as translation operations. This method allows TransNet to capture richer semantic information, enhancing the Social Relation Extraction task and providing deeper insights into vertex relationships.
A novel sentiment-guided review-aware recommendation model (SentiRec) (Hyun et al., 2018) efficiently processes review data by transforming reviews into sentiment-encoded vectors. This significantly reduces computational demands and captures essential sentiment details, enabling tailored recommendations based on emotional content. Additionally, the Word-Driven and Context-Aware Review Modeling for Recommendation (WCN-MF) (Wang et al., 2018) model enhances text analysis by integrating Deep Latent Dirichlet Allocation with CNNs, allowing for a detailed extraction of contextual features from reviews, which traditional pooling methods might miss. This method improves the model’s contextual understanding and recommendation relevance.
The Weighted Aspect-Based Opinion Mining Using Deep Learning for Recommender Systems (AODR) (Da’u et al., 2020) focuses on extracting weighted, aspect-based opinions, effectively identifying the relative importance users assign to different product aspects, thereby aligning recommendations more closely with user preferences. A hybrid Neural network to Combine Textual information and Rating for recommendation( NCTR ) (Liu et al., 2021a) integrates CNNs with collaborative filtering to address data sparsity, uncovering latent user-item interactions, and enriching the model’s understanding of complex preference patterns. Finally, Deep Rating and Review Neural Network for Item Recommendation (DRRNN) (Xi et al., 2021) leverages both ratings and reviews during training, ensures a comprehensive preservation of semantic information, which enhances personalization and offers a more holistic understanding of user and item dynamics.
Together, these models demonstrate the flexibility of CNNs in adapting to different aspects of recommendation challenges, from general feature extraction to coarse-grained and fine-grained sentiment analysis, highlighting their broad applicability across diverse recommendation scenarios.
Neural Attention-based Approaches: Attention mechanisms within recommender systems significantly advance traditional algorithms by focusing on the most relevant parts of the data. This method effectively differentiates crucial information from noise, which enhances the interpretability and relevance of the recommendations. Broadly, these models excel in managing the detailed information present in user reviews, which significantly boosts the specificity and personalization of recommendations.
The Adaptive Aspect Attention Model for Rating Prediction (A3NCF) (Cheng et al., 2018a) uses an aspect-based attention framework that selectively focuses on various aspects of items as valued by users, extracting preferences and item characteristics from reviews. This offers a refined and aspect-oriented recommendation model. This specificity is beneficial in tailoring recommendations, though A3NCF may face challenges when reviews are sparse or superficial, affecting its ability to fully capture user preferences. Similarly, the Attentive Interaction Network (AIN) (Mei et al., 2018) extends these capabilities to Context-Aware Recommender Systems (CARS) (Kulkarni and Rodd, 2020) with a multi-module approach that includes interaction-centric, user-centric, and item-centric components. This comprehensive framework allows AIN to dynamically adjust the contextual relationships between users and items, providing deeply personalized recommendations tailored to current user situations. However, AIN might struggle in environments where contextual information is either too generic or excessively varied.
The Attentive Aspect Modeling for Review-aware Recommendation (AARM) (Guan et al., 2019) specifically tackles the aspect-sparsity issue prevalent in many reviews. By learning from both consistent and varying aspect connections between users and products, AARM significantly enhances the depth of understanding in user-item relationships. Meanwhile, Cross-Modality Mutual Attention (NRCMA), (Luo et al., 2021) utilizes a two-tower structure to synchronize user and item representations through mutual attention, effectively enhancing the match between user preferences and item characteristics, though it faces challenges in very dynamic environments.
Lastly, the Enhanced review-based rating prediction (ERP) (Wu et al., 2021) incorporates a mechanism to evaluate the general influence of reviewers, allowing it to differentiate and tailor recommendations based on how much trust to place in each review. This capability ensures that the recommendations are not only personalized but are also aligned with the most credible and relevant user feedback, enhancing the accuracy and relevance of the suggestions.
Together, these attention-based models enhance recommender systems by improving how user data is interpreted and utilized. While all models enhance traditional methods by incorporating advanced attention mechanisms, they vary in applications and challenges. A3NCF focuses on aspect-based ratings, AIN adapts to contextual variations, AARM addresses review aspect sparsity, and NRCMA bridges user-item representation gaps.
Integrated CNN and Attention-based Approaches: This approach combines the strengths of Convolutional Neural Networks (CNNs) and attention mechanisms to enhance the accuracy of the recommender systems. By utilizing CNNs, this approach effectively extracts and learns deep, complex features from textual data, such as user reviews or item descriptions. The incorporation of attention mechanisms further refines the process by selectively focusing on the most relevant parts of the data, ensuring that the most informative aspects are emphasized. This integrated approach enables a more detailed understanding of user preferences and item characteristics, leading to more accurate and personalized recommendations.
The Interpretable Convolutional Neural Networks with Dual Local and Global Attention (D-Attn) (Seo et al., 2017) utilizes a sophisticated dual attention mechanism that enables a detailed analysis of textual data. By merging local and global attention, D-Attn not only highlights crucial phrases and contextual indicators within reviews but also provides a more comprehensive understanding of the content. This dual-layered attention significantly enhances interpretability over traditional models like Matrix Factorization, HFT (McAuley and Leskovec, 2013), and Convolutional Matrix Factorization, making it proficient in managing complex datasets where detailed textual analysis is critical. In contrast, the Neural Attentional Rating Regression with Review-level Explanations (NARRE) (Chen et al., 2018a) focuses on the qualitative value of reviews rather than just numerical ratings. Its unique attention mechanism assesses the usefulness of each review, providing detailed explanations that enhance user trust and deepen satisfaction. NARRE proves to be particularly effective in environments that depend heavily on rich user feedback for recommendation refinement, making it ideal for platforms that prioritize depth and explanation in user interactions.
Context-Aware User-Item Representation Learning (CARL) (Wu et al., 2019a) and Dual Attention Mutual Learning (DAML) (Liu et al., 2019a) both excel in dynamically adapting recommendations based on user and item interactions. CARL leverages CNNs enhanced with an attention layer to derive insights from both review and interaction data, enriching the contextual understanding necessary for precise personalization. Demonstrating this, DAML integrates local and mutual attention mechanisms, which balance the representation of users and items to ensure accuracy and relevance in its outputs. Neural Recommendation with Personalized Attention (NRPA) (Liu et al., 2019b) personalizes recommendations by focusing on the informativeness of reviews for different users and items. The Aspect-aware Neural Review Rating (ARP) (Wu et al., 2019b) focuses on aspect-level opinions in reviews, providing a detailed understanding of user preferences and item qualities. By using a collaborative learning framework to train both review-level and aspect-level rating predictors, ARP improves rating prediction accuracy, particularly for items with sparse ratings.
Multilingual Review-aware Deep Recommender System (MrRec) (Liu et al., 2021d) tackles the complexities of handling reviews in multiple languages by integrating an aspect-based sentiment analysis module with a multilingual recommendation framework. This model efficiently processes and leverages fine-grained user-item interactions from diverse linguistic backgrounds. To effectively capture the semantic features of reviews, a CNN is used to perform convolution operations on each language and aspect-specific matrix. Additionally, attention mechanisms are utilized to emphasize significant features and interactions. Meanwhile, RPRM (Wang et al., 2021b) utilizes review properties to assess their usefulness, introducing two new loss functions and a negative sampling strategy that directly link these properties with user preferences. This method allows RPRM to provide a more refined modeling of user preferences and item attributes.
Overall, the integration of CNNs with attention mechanisms in recommender systems enhances the precision and personalization of recommendations by effectively analyzing complex textual data.
Recurrent Neural Networks (RNNs)-based Approaches: RNN and their variants, such as LSTM, and GRUs, are particularly effective in review-based recommender systems. By leveraging their ability to process sequential information, these models can analyze the detailed sentiment and progression within user reviews. This facilitates a deeper understanding of preferences and item characteristics from textual feedback. Consequently, RNNs enable recommender systems to generate more accurate and personalized recommendations by addressing complex details in user feedback.
The Sentiment-aware Deep Recommender System with neural Attention network (SDRA) (Da’u and Salim, 2019) further enhances the recommendations by fully leveraging user-generated textual reviews. Integrating a semi-supervised topic model with a deep learning framework (e.g. LSTM) through a neural attention mechanism, it adeptly captures both product aspects and the associated user sentiments from textual reviews, leading to improved recommendation performance. The use of Long Short-Term Memory (LSTM) encoders for capturing contextual word information and a co-attention mechanism for fine-tuning user-item interactions based on these insights further enhances performance. SDRA offers a pathway for incorporating user sentiment and aspect information into the recommendation process, showcasing its potential for more nuanced and effective recommendations. The Topical Attention Regularized Matrix Factorization (TARMF) (Lu et al., 2018) integrates attention-based GRUs with matrix factorization, tackling the inherent sparsity in user-item ratings that frequently challenges collaborative filtering systems. TARMF’s hybrid approach not only enhances sentiment analysis of product features but also adds a layer of transparency and explainability, setting it apart from traditional matrix factorization methods. The Selective Distillation Network (SDNet) (Chen et al., 2018b) introduces a distillation approach where the system uses external knowledge, such as user reviews, to train a “teacher” model that guides a more streamlined “student” model that focuses on rating prediction from user-item pair. The “teacher” model uses both CNN and LSTM for review processing to extract the enriched features from review. This separation allows for efficient runtime operation without sacrificing the depth of knowledge extraction, addressing the dual needs of performance and efficiency in dynamic environments.
Leveraging Ratings and Reviews with Gating Mechanism for Recommendation (RRGM) (Xia et al., 2019) and Dynamic Explainable Recommender (DER) (Chen et al., 2019) both emphasize the dynamic nature of user preferences. While the former uses dual attention-based GRU networks to adapt recommendations based on current user-review interactions, DER employs a time-aware GRU to track evolving preferences, supplemented by a CNN for item profiling. These models excel in environments where user preferences are continually changing, offering recommendations that are not only timely but also explainable. Aspect-based Fashion Recommendation with Attention Mechanism (AFRAM) (Li and Xu, 2020), focusing on the fashion industry, processes both local and global aspect features to provide precise fashion advice, demonstrating its capability to enhance recommendation accuracy in a specialized market with unique consumer demands.
Lastly, Asymmetrical Hierarchical Networks (AHN) (Dong et al., 2020) and Reliable Recommendation with Review-level Explanations (RRRE) (Lyu et al., 2021) focus on the granularity and reliability of reviews respectively. AHN addresses the varying scope of user and item reviews by applying differently focused attention modules, enhancing the relevancy of content extracted for recommendations.Specifically, the model utilizes bi-directional LSTM networks to process the sequential data within reviews, allowing for the effective capture and integration of contextual and temporal relationships between sentences. Conversely, RRRE prioritizes the authenticity of reviews, using mechanisms to assess and filter content based on reliability, which is crucial for maintaining the integrity of recommendation outputs. Overall, these RNN-based models improve recommendation accuracy and user satisfaction by processing sequential review data. They vary in their specific applications—some enhance interpretability and trust (TARMF and RRRE), while others focus on adapting to changing user conditions (DER and RRGM). The use of RNNs alongside technologies like attention mechanisms and adversarial networks indicates the potential for ongoing advancements in the field, aiming to align recommender systems more closely with user expectations and preferences.
Graph Neural Network-based Approaches: Graph Neural Networks (GNNs) have significantly advanced in dealing with complex user-item interactions in recommendation systems. By exploiting the structural connectivity of data, GNNs excel in capturing high-order relationships. This is particularly effective for addressing data sparsity, cold start problems, and improving the interpretability of recommendations through detailed insights from user and item interactions.
Tripartite Graph Ranking (TrinRank) (He et al., 2015) evolves beyond traditional collaborative filtering by modeling ternary relationships in a heterogeneous tripartite graph (user-item-aspect). The model focuses on vertex ranking in the tripartite graphs, tailored for personalized recommendation which provides explainability and transparency by modeling aspects in reviews. In contrast, Review Meet Graphs (RMG) (Wu et al., 2019c) utilizes a multi-view framework that merges review texts with user-item graphs within a GNN architecture. By doing so, RMG captures detailed user-item interactions more adeptly, highlighting the advantage of integrating diverse data views to boost the effectiveness of recommender systems.
Heterogeneous Graph Neural Recommender (HGNR) (Liu et al., 2020) and Set-Sequence-Graph (SSG) (Gao et al., 2020b) both address the fundamental challenges of recommender systems. HGNR focuses on alleviating the cold start problem by using social information and reviews to predict links between users and items, demonstrating GNNs’ capability to leverage heterogeneous data. Conversely, SSG enhances recommendation accuracy by combining sequential and graphical views, which provides a comprehensive understanding of short-term user preferences and broader collaborative signals.
Review Graph Neural Network (RGNN) (Liu et al., 2020) tackles bias and noise in data by constructing individual review graphs for each user/item pair. This model emphasizes the importance of understanding both local and global word dependencies in reviews, thus enhancing the precision of rating predictions through a sophisticated graph attention mechanism and personalized pooling strategies. Finally, Aspect-Aware Higher-Order Representations (AHOR) (Wang et al., 2023a) utilize heterogeneous graphs to capture aspect-aware relationships, enriching the learning of high-order user and item representations through multi-hop connectivity. The incorporation of an attention mechanism to selectively focus on relevant features ensures that AHOR not only enhances performance but also significantly improves the interpretability of recommendations.
Overall, these GNN-based models utilize graph structures to analyze relationships but vary in their approach—some focus on broad interaction patterns, while others concentrate on detailed aspects of content and context. This variety highlights the potential for developing hybrid models that could merge these approaches to enhance accuracy and customization in recommendations.
Contrastive Learning-based Approaches: The Contrastive Learning-based Approach (Khosla et al., 2020; Xia et al., 2022) refines recommender systems by its unique power of data representation learning technique - modeling the data representation by minimizing the distance between similar data points while maximizing the distance between similar data points. This approach has been successful in tasks requiring detailed data understanding, and making recommendations based on user reviews.
The Review-aware Graph Contrastive Learning (RGCL) framework leads this approach by enhancing user and item embeddings through the integration of review data. It utilizes contrastive learning techniques to improve node embedding and interaction modeling, thus addressing data sparsity and more accurately capturing user preferences. Similarly, the Review-based Multi-intention Contrastive Learning (RMCL) (Yang et al., 2023) method introduces learning intention representations from reviews based on a mixed Gaussian distribution hypothesis. It uses a multi-intention contrastive learning strategy to link user and item reviews, fostering a deeper connection between them.
The Multi-Aspect Graph Contrastive Learning (MAGCL) (Wang et al., 2023b) advances the application of contrastive learning in GNNs by incorporating multi-aspect graph representations and multi-task learning. This model breaks down user-generated reviews into distinct semantic aspects, enhancing the granularity of user preference analysis. Its contrastive learning module not only improves the model’s predictive accuracy but also helps to overcome the common over-smoothing problem in GNNs. Recently, Multi-Level Cross-modal Contrastive Learning (MCCL) (Wei et al., 2024) explores contrastive learning across different modalities, enhancing the ability of models to leverage semantic information in a self-supervised way.
Overall, these models utilize contrastive learning to refine how recommender systems process review data using diverse and specific methods and different complexities. RGCL and RMCL focus on enhancing node and intention understanding, respectively, while MAGCL and MCCL expand the scope to multi-aspect and cross-modal analysis.
LLMs-based Approaches: Large Language Models (LLMs) excel at processing complex text, which enhances the personalization and clarity of recommender systems (Laskar et al., 2023a, b). A significant advancement in this area has been achieved by integrating LLMs, specifically GPT-2, into the recommendation process (Li et al., 2023b). By utilizing a prompt tuning technique, the model dynamically extracts key aspects reflecting user preferences and item properties directly from textual reviews. These extracted aspects are then used to inform and personalize the recommendation process, leading to a more tailored and insightful user experience.
This integration not only improves the relevance and quality of recommendations but also contributes to the transparency and explainability of the system, addressing a critical need in recommender systems research. Evaluated on three large-scale datasets, the model demonstrates superior performance over existing methods, showcasing its potential for practical application and scalability. This study demonstrates the transformative potential of LLMs in recommender systems, paving the way for further exploration and development in this exciting intersection of natural language processing and personalized recommendation.
Table 3. The detailed summary of the Deep Learning-based Approaches
Category Models Year Architecture Features Datasets Evaluation Tasks
ConvMF(Kim et al., 2016) 2016 CNN, MF reviews amazon, yelp & beer RMSE accuracy & interpretability
DeepCoNN (Zheng et al., 2017) 2017 CNN reviews yelp MSE accuracy
TransNet(Catherine and Cohen, 2017) 2017 CNN reviews yelp MSE accuracy
JRL(Chen et al., 2018b) 2017 MLP rat, img & rev amazon NDCG accuracy
D-Attn(Seo et al., 2017) 2017 CNN, Attn reviews yelp & amazon MSE accuracy & interpretability
NARRE(Chen et al., 2018a) 2018 CNN, Attn reviews yelp & amazon MSE accuracy
SentiRec(Hyun et al., 2018) 2018 CNN reviews amazon MSE scalability & memory efficient
WCN-MF(Wang et al., 2018) 2018 CNN, LDA, MF reviews amazon RMSE accuracy
A3NCF(Cheng et al., 2018a) 2018 Attn ratings & reviews amazon & yelp MSE accuracy
TARMF(Lu et al., 2018) 2018 GRU, Attn ratings & reviews amazon & yelp MSE accuracy
AIN(Mei et al., 2018) 2018 Attn ratings & reviews yelp, food & frappe RMSE accuracy & interpretability
ANR(Chin et al., 2018) 2018 Attn reviews yelp & amazon MSE, MAE accuracy & interpretability
SDNet(Chen et al., 2018b) 2018 CNN, LSTM review amazon MSE accuracy
DL CARL(Wu et al., 2019a) 2019 CNN, Attn reviews amazon MSE accuracy
NRPA(Liu et al., 2019b) 2019 CNN reviews amazon MSE accuracy
DAML(Liu et al., 2019a) 2019 CNN, MF reviews amazon MAE accuracy & interpretability
AARM(Guan et al., 2019) 2019 Attn reviews yelp & amazon N & HR accuracy & interpretability
ARP(Wu et al., 2019b) 2019 Attn ratings & reviews chines res. reviews MAE accuracy
SDRA(Da’u and Salim, 2019) 2019 LSTM reviews amazon & yelp MSE accuracy
ATR(Rafailidis and Crestani, 2019) 2019 adversarial net reviews amazon MSE accuracy
RACRec(Jin et al., 2020) 2020 CNN reviews amazon RMSE accuracy
RSBM(Huang et al., 2020b) 2020 CNN, Attn reviews amazon & yelp MSE accuracy
AFRAM(Li and Xu, 2020) 2020 CNN, LSTM reviews amazon MSE, MAE accuracy
AODR(Da’u et al., 2020) 2020 CNN reviews amazon MSE, RMSE accuracy
SSG(Gao et al., 2020b) 2020 CNN review amazon & yelp MAE accuracy
NCTR(Liu et al., 2021a) 2021 CNN reviews & ratings amazon & movie MSE, MAE accuracy
RPRM(Wang et al., 2021b) 2021 CNN reviews amazon P, R & M accuracy
NRCMA(Luo et al., 2021) 2021 CNN reviews amazon MSE accuracy
Acapnet(Yang et al., 2021) 2021 CNN reviews amazon MSE accuracy
MrRec(Liu et al., 2021d) 2021 CNN, Attn reviews amazon & goodreads MSE acc, diversity & cold start
SIFN(Zhang et al., 2021) 2021 Attn review amazon MSE accuracy
DRRNN(Xi et al., 2021) 2021 CNN, MLP reviews & ratings amazon RMSE data sparsity
U-ARM(Sun et al., 2021) 2021 DNN reviews amazon & yelp RMSE accuracy & interpretability
ERP(Wu et al., 2021) 2021 CNN reviews & ratings amazon & yelp MSE, RMSE accuracy
HACN(Wu et al., 2021) 2021 CNN, Attn reviews amazon MAE, RMSE accuracy
RGCL(Shuai et al., 2022) 2022 GNN, CL reviews amazon & yelp MSE accuracy
RPR(Liu et al., 2022a) 2022 CNN reviews amazon & yelp MSE & MAE accuracy
AHOR(Wang et al., 2023a) 2023 GNN reviews amazon & yelp NDCG, HR accuracy
MCCL(Wei et al., 2024) 2024 GNN & CL reviews amazon & yelp NDCG, P accuracy
Reinforcement Learning-based Approaches: In Reinforcement Learning (RL), an agent learns to make decisions by interacting with an environment (Wiering and Van Otterlo, 2012). The agent receives rewards or penalties based on its actions, which guide it toward optimizing its behavior over time. In review-based recommender systems (RBRS), RL can be particularly effective in scenarios where user preferences are dynamic and the system must adapt to continuously evolving data. RL algorithms can optimize the recommendation process by learning from user interactions and feedback, such as reviews, to enhance the personalization and accuracy of the recommendations (Afsar et al., 2022).
A recent paper explores the use of RL for generating shilling attacks in RBRS, highlighting a novel vulnerability where fake reviews are used to manipulate recommendation outcomes. The Shilling Black-box Review-based Recommender Systems (Chiang et al., 2023), utilize adversarial training to test and enhance system robustness against fake reviews that manipulate recommendation outcomes. This approach not only demonstrates the system’s susceptibility to shilling attacks but also its capacity to defend against such tactics, emphasizing the need for robust security measures in recommender systems.
This research opens up areas for exploring reinforcement learning in review-based recommender systems. Future work could focus on developing more sophisticated RL algorithms to enhance the system’s resilience. Additionally, exploring the integration of RL with other machine learning techniques could lead to more robust and adaptable recommender systems that can handle increasingly complex user interactions and preferences.
Miscellaneous Approaches: The Sentiment Utility Logistic Model (SULM) (Bauman et al., 2017) utilizes sentiment analysis to identify key aspects of products that influence user preferences, enhancing the personalization of recommendations. For example, it might recommend a restaurant specifically for its standout seafood, effectively tailoring suggestions to user preferences. Demonstrated in various settings like restaurants and hotels, SULM’s aspect-focused recommendations have been well-received for improving user satisfaction. Building on this focus on user insights, the Adversarial Training for Review-based Recommendations (ATR) (Rafailidis and Crestani, 2019) model integrates user reviews with collaborative filtering. By using adversarial training, this model aligns deep representations of reviews with user and item features, aiming to improve the accuracy of recommendations.
The Sentiment-aware Interactive Fusion Network (SIFN) (Zhang et al., 2021) addresses the precision of sentiment interpretation in recommendations by utilizing Bidirectional Encoder Representation From Transformer (BERT) (Devlin et al., 2018) for encoding and a lightweight sentiment learner for semantic analysis. This model enhances the accuracy of item suggestions by using explicit sentiment labels, although it faces challenges with sparse or vague reviews. Counterfactual Explainable Recommendation (CountER) (Tan et al., 2021) leverages counterfactual reasoning to provide clear and meaningful explanations for recommendations by identifying minimal changes in item aspects that could influence user decisions. CountER not only enhances the transparency of recommendations but also assesses the impact and clarity of its explanations, significantly outperforming other methods in explainability.
The Lightweight Unbiased Multi-teacher Ensemble (LUME) (Xv et al., 2022) addresses computational complexity, biased recommendations, and poor generalization by implementing a multi-teacher ensemble with debiased knowledge distillation. This strategy aggregates insights from various pretrained models to form a streamlined, unbiased system that excels in generalization. The detailed summary of the deep learning-based approaches are presented in Table 3.
6. Datasets
Traditional rating-based recommender systems rely on the dataset that comes with user, item, and numerical rating information, e.g. MovieLens Dataset  [^4] https://grouplens.org/datasets/movielens/, IMDb Dataset [^5] https://developer.imdb.com/non-commercial-datasets/, and Goodreads Dataset [^6] https://www.goodreads.com/genres/dataset. In contrast, review-based recommender models are evaluated mainly with the datasets that contain user, item, and users’ textual review feedback, e.g. Amazon [^7] https://jmcauley.ucsd.edu/data/amazon/, Yelp [^8] https://www.yelp.com/dataset, TripAdvisor [^9] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/, and Beer reviews  [^10] https://cseweb.ucsd.edu/~jmcauley/datasets.html#multi_aspect.
  • Amazon Product Review: This is one of the largest and most commonly used product review datasets, originating from Amazon website. Spanning from May 1996 to July 2014, it contains over 143 million reviews. With 21 distinct categories of items, this dataset is the most extensive publicly accessible product review dataset to date, encompassing user ID, item ID, rating, and textual feedback.
  • Yelp Review: It is the largest restaurant review dataset containing more than 1M reviews and ratings. The dataset is collected from Yelp.com, an online platform for restaurant reviews. This dataset also comes with user, and item IDs along with textual reviews.
  • Beer Review: The Beer review dataset is also often used to evaluate the recommendation model. It is collected from the ratebeer.com website. Data was extracted for a period of 10 years, up to November 2011, comprising approximately 3 million reviews
  • TripAdvisor Review: Tripadvisor.com is a tourism management platform, which provides a range of services such as hotel bookings, flight reservations, restaurant recommendations, cruise options, and car rental facilities. The data was collected over an extensive 8-year period, spanning from 2004 to 2012. Within this dataset, one can find comprehensive evaluations including overall ratings, multi-criteria assessments, and textual reviews. There exist a total of 8 distinct criteria, each accompanied by its respective rating: value, rooms, location, cleanliness, check-in/front desk, service, sleep quality, and business service. Ratings for both individual criteria and overall experiences fall within a scale from -1 to 5.
The detailed statistics of datasets including each category of Amazon review, and Yelp datasets are presented in Table 4.
Table 4. Data statistics
Datasets Number of users Number of items Number of reviews Density
Cell Phones 27879 10429 194439 0.067%
Digital Music 5,541 3,568 64,706 0.330%
Home and Kitchen 66519 28237 551682 0.03%
Foods 14,681 8,713 151,254 0.118%
Automotive 2,928 1,835 20,473 0.99%
Baby 19445 7050 160792 0.122%
Office Product 4,905 2,420 53,258 0.449%
Toys and games 19,412 11,924 167,597 0.072%
Instant Video 5,130 1,685 37,126 0.433%
Musical Instruments 1,429 900 10,261 0.798%
Video games 24,303 10,672 231,577 0.089%
Beauty 22,363 12,101 198,502 0.0734%
Clothing, Shoes, and Jewelry 39,387 23,033 278,677 0.0307%
Movies and TV 123,960 50,052 1,697,533 0.0274%
Sports Outdoors 35,598 18,357 296,337 0.045%
Office Products 4,905 2,420 53,228 0.449%
Pet Supplies 18,070 8,508 155,692 0.99%
CDs and Vinyl 75,258 64,443 1,097,494 0.023%
Tools and Home Improvement 15,438 10,217 133,414 0.99%
Grocery & Gourmet Food 127487 41320 1143470 0.22%
Electronics 2,832 19,816 53,295 0.095%
TripAdvisor Hotel 429,928 3,828 620,172 0.9996%
Yelp 45,981 11,537 229,907 0.0433%
Beer 40,213 110,419 2,924,127 0.05%
7. Evaluation Metrics
Research on review-based recommender systems focuses on two core tasks: rating prediction and top-N items recommendation. For the rating prediction task, the problem is formulated as a regression problem and the models are evaluated with MSE, RMSE, and MAE. In contrast, top-N items recommendation tasks are formulated as a ranking problem, evaluating them with NDCG@N, recall@N, precision@N, F1@N, HR, MRR, and MAP. While the former helps to evaluate the user’s sentiment prediction accuracy, the latter explores the relevancy of the list of recommended items. Each of the evaluation metrics is defined below.
7.1. Rating Prediction
For a given user $u$ and an item $i$, the rating prediction task involves estimating the overall sentiment or overall ratings ($\hat{r}$), indicating the user $u^{\prime}s$ overall preferences towards the item $i$. If the true rating is $r$, the Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) are defined as follows,
(6) $MSE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{r_{i}}-r_{i})^{2}}$
(7) $RMSE=\sqrt{\frac{\sum_{i=1}^{N}(\hat{r_{i}}-r_{i})^{2}}{N}}$
and
(8) $MAE=\frac{1}{N}\sum_{i=1}^{N}|\hat{r_{i}}-r_{i}|$
where $N$ is the total number of data points used for testing.
7.2. Top-N items Recommendation
The objective of the top-N items recommendation task is to recommend top-N most relevant items to users. It is important to recommend the most relevant items and maintain the ranking quality.
NDCG (Normalized Discounted Cumulative Gain): NDCG (Normalized Discounted Cumulative Gain) is a ranking quality metric, which measures the relevancy and the ranking quality of the recommended items. For top-N recommended items, NDCG is calculated as
(9) $NDCG@N=\frac{DCG@N}{iDCG@N}$
where $DCG$ measures the Discounted Cumulative Gain for the recommended order whereas iDCG measures the ideal order. The NDCG depends on the two terms: CG (Cumulative Gain ) and DCG (Discounted Cumulative Gain). CG is defined as the sum of the relevance score for all the recommended items, and Discounted Cumulative Gain takes into account the positions of the relevant items and they are computed as
(10) $CG=\sum_{i=1}^{N}relenaceScore_{i}$
and
(11) $DCG=\sum_{i=1}^{N}\frac{relenaceScore_{i}}{\log_{2}(i+1)}$
MRR (Mean Reciprocal Rank): MRR (Mean Reciprocal Rank) focuses on retrieving the first relevant item from the top-N recommended items.
(12) $MRR=\frac{1}{U}\sum_{u=1}^{U}\frac{1}{rank_{u}}$
where $U$ is the total number of users and $rank_{u}$ is the rank of the first relevant item for user u.
HR (Hit Ratio): Hit Ratio is one of the most widely used top-N item recommendation evaluation metrics. HR computes the share of users that receive at least one relevant item from the recommended top-N items. HR is computed as
(13) $HR=\frac{U_{hit}^{L}}{U}$
where $U$ is the total number of users and $U_{hit}^{L}$ is the number of users with at least one relevant recommendation.
MAP (Mean Average Precision): MAP is another popular top-N item recommendation evaluation metric, which measures the average precision across all relevant ranks within the top-N recommended items in the list. It is computed as follows,
(14) $MAP=\frac{1}{|U|}\sum_{u=1}^{|U|}AP(u)$
where AP(u) is the average precision for the user $u$. Average Precision (AP) at postiion k can be defined as
[…]
where $m$ is the total number of relevant items for a particular user. Precision(k) is the precision for each position and rel(k) represents 1 if the item at position k is relevant otherwise it is 0.
Precision: Precision measures how many of the recommended items are relevant. For Top-N recommended items, the Precision@N is defined as
(16) $Precision@N=\frac{\text{Number of relevant items in the Top N}}{N}$
Precision for top-N recommendation problems depends on the number of recommended items.
Recall@N: Recall@N represents out of the total relevant items how many relevant items can be recommended from top-N items. It is defined as
(17) $Recall@N=\frac{\text{Number of relevant items in the Top N}}{\text{Number of relevant items}}$
F1@N: F1 measures the balance between precision and recall at N, measuring the harmonic mean of precision and recall, and is computed as
(18) $F1@N=2*\frac{\text{Precision@N}\times\textit{recall@N}}{\text{Precision@N + recall@N}}$
8. MAJOR CHALLENGES IN REVIEW-BASED RECOMMENDER SYSTEMS
Review-based recommender systems uniquely incorporate user-generated content to predict preferences and suggest items. While these systems leverage detailed insights from textual reviews, they encounter specific challenges. This section outlines these unique challenges, discussing both historical solutions and areas where challenges persist, emphasizing why these issues continue to impact the effectiveness of Review-based recommender systems.
8.1. Representation Learning
A primary challenge in developing effective review-based recommender systems lies in learning meaningful representations of user behavior and item attributes from reviews. Users’ textual reviews are unstructured and inherently noisy and often contain slang and emojis. Extracting meaningful information from such complex data is challenging (Almahairi et al., 2015; Baral and Li, 2016; Cheng et al., 2018a). Various deep learning and NLP techniques have been utilized in existing approaches to model the complex nature of unstructured text data. Early approaches relied on Bag of Words (BoW) and word embedding methods to extract features from reviews. However, these approaches struggled to learn contextual representations from text, leading to sub-optimal recommendation performance. Subsequently, Convolutional Neural Networks (CNNs) were introduced to learn contextual representations from reviews. While CNN-based methods proved successful, they often failed to capture users’ detailed preferences from reviews. It became evident that certain aspects of reviews provide more valuable information than others, highlighting the need to learn dynamic preferences. To address this, attention mechanisms were found to be effective. By utilizing attention mechanisms with CNN or RNN, relevant parts of reviews are captured, thereby enhancing the recommendation performance as well the interpretability. Very recently, Contrastive Learning (Khosla et al., 2020) in conjunction with graph neural networks has demonstrated the superiority of the existing classical deep learning method in terms of user/item representation learning as well as the recommendation accuracy and interpretability.
8.2. Incorporating Reviews into Recommender Systems
The key to enhancing the performance of review-based recommender systems lies in effectively incorporating review features into the recommendation process. Reviews are integrated in two main ways: 1) by creating user/item profiles from accumulated reviews to enrich user and item representation, and 2) by combining review features with rating features. However, simply aggregating reviews to construct user or item profiles can overlook the dynamic nature of user preferences, which are influenced by contextual factors like time, location, mood, and trends. Traditional methods Bag of Words ( e.g. GloVe Word Embedding (Pennington et al., 2014) or Word2Vec (Church, 2017) ), use a fixed-size vocabulary and may not capture these details, leading to potential inaccuracies in profiling. Moreover, when merging review features with ratings, there is a risk of mismatch—sentiments expressed in reviews might differ from the ratings given. Addressing these challenges requires more sophisticated models that account for the temporal and contextual dynamics of user decisions, thereby enhancing both the accuracy and interpretability of the recommendations.
8.3. Scalability
Handling the vast volumes of review data on e-commerce platforms introduces significant scalability challenges. These include managing sparse and sometimes unclear reviews, extracting meaningful insights efficiently, and accommodating the sheer volume of data processed. The scalability of recommender systems is further complicated by the need to adapt to the dynamic nature of user preferences and item popularity, which may shift due to factors like seasonal demand, marketing efforts, or new product launches. Such systems must be capable of updating in real-time to reflect these changes, ensuring recommendations remain relevant and effective (Xiong et al., 2021).
The integration of diverse data types further amplifies scalability challenges (Choi et al., 2022). Modern e-commerce platforms do not rely solely on textual reviews; they also incorporate images, videos, and audio feedback, creating a multimodal dataset. Each data type requires specific processing techniques to extract meaningful features that can inform the recommendation process. For instance, NLP techniques are suited for textual data, while image processing algorithms are necessary for analyzing user-uploaded photos. Successfully merging these varied data sources demands sophisticated data fusion techniques and machine learning models that can handle multimodality without compromising the efficiency or scalability of the recommender system.
8.4. Privacy and Ethical Considerations
As review-based recommender systems increasingly rely on user-generated content to enhance personalization and accuracy, addressing privacy, ethical handling, and bias becomes crucial. User reviews often contain personal identifiers and data that, if mismanaged or exposed, could result in privacy breaches. Additionally, text in reviews can naturally carry biases from the user, introducing potential discrimination into the system. Tackling these biases, detecting them, and mitigating their influence within recommendation algorithms are vital challenges. Ethical considerations around data collection, consent, and usage are also paramount to maintain user trust and regulatory compliance. Developing systems that prioritize user privacy through secure data handling, anonymization techniques, and transparent user consent processes is essential. Moreover, creating algorithms that prevent discrimination and ensure fairness in recommendations is crucial to foster an inclusive and trustworthy digital environment.
9. DISCUSSION AND FUTURE DIRECTIONS
The study of review-based recommender systems highlights both the progress and challenges in this area. Despite encountering several challenges, review-based recommender systems offer a significant advancement in recommendation technology by leveraging the fine-grained insights present in user-generated reviews to provide highly personalized suggestions. These systems explore the textual content of reviews to grasp users’ sentiments, preferences, and fine-grained opinions about products or services. This deeper analysis facilitates enhanced personalization, as the systems can comprehend individual tastes and requirements with greater precision, resulting in recommendations that closely align with each user’s specific preferences. Moreover, review-based systems improve accuracy by considering qualitative information present in reviews, such as detailed explanations, experiences, and comparisons, which may not be captured by ratings.
Additionally, review-based recommender systems excel at capturing detailed information about products or services, providing recommendations that consider specific attributes or characteristics that matter most to users. This adaptability to diverse preferences ensures that the recommendations can meet various user interests, accommodating different priorities such as quality, price, or specific features. By leveraging user-generated content, review-based systems offer a trustworthy recommendation process, fostering confidence in users’ decision-making. However, these improvements also bring up important questions about ethics, how transparent the algorithms are, and the importance of maintaining user trust and data security.
Given the complexities and rapid advancements in Review-based Recommender System, the following subsection outlines several promising areas for future research in this field:
Table 5. Future Directions in Review-based Recommender Systems
Area of Focus Proposed Directions Impact
Enhancing Interpretability and Trust Develop models linking recommendations to aspects in reviews; incorporate explanation modules. Improves user trust and satisfaction by making recommendation logic transparent and understandable.
Leveraging Multimodal Data Integrate text with images, videos, and audio to enrich user profiles and item features. Enhances depth of insights into user preferences and improves recommendation accuracy.
Leveraging Multi-criteria Rating Data Combine fine-grained aspect ratings with reviews to refine recommendations. Enhances performance and interpretability, addressing data sparsity issues.
Addressing Data Sparsity and Cold Start Use user demographics, item metadata, and transfer learning to improve new user/item recommendations. Mitigates cold start problems and enhances recommendation accuracy for new users/items.
Ethical Considerations and Bias Mitigation Develop mechanisms to detect and mitigate biases; ensure fairness and diversity. Builds trust and compliance, ensuring recommendations are equitable and unbiased.
Real-time Adaptation and Personalization Implement dynamic models that adjust recommendations based on real-time data. Keeps recommendations relevant and responsive to current user behavior and trends.
Cross-Domain Recommendations Apply domain adaptation techniques to leverage knowledge across product areas. Improves recommendation flexibility and consistency across various product categories.
Large Language Models and Recommendations Use LLMs for analyzing reviews and generating relevant suggestions. Leverages advanced NLP capabilities to enhance recommendation quality and relevance.
Responsible Recommendations Focus on privacy, diversity, and fairness in system design and operation. Ensures that recommendation systems operate responsibly, promoting user trust and societal benefit.
9.1. Enhancing Interpretability and Trust
Future advancements should focus on the interpretability and trust. Users benefit not only from accurate recommendations but also from understanding why specific items are recommended. Enhancing interpretability involves developing models that can explicitly link recommendations to specific aspects mentioned in reviews, such as quality, usability, or aesthetics. This can be achieved through more sophisticated aspect-based analysis and by incorporating explanation modules that can convey the rationale behind recommendations in an easily understandable manner. Building systems that offer transparent reasoning behind their suggestions could significantly improve user trust and satisfaction.
9.2. Leveraging Multimodal Data
As user reviews increasingly include multimedia content in reviews, such as images, videos, and audio clips, future recommendation systems could benefit from a multimodal approach that integrates these diverse types of data. This involves not only analyzing textual content but also understanding and extracting sentiments and preferences from images and videos. For instance, visual sentiment analysis from product images or user-generated content in reviews could offer additional insights into user preferences, enhancing the recommendation process. Research in multimodal learning and its integration into recommender systems remains an under-exploited area with considerable potential for innovation.
9.3. Leveraging Multi-criteria Rating Data
Incorporating multi-criteria ratings, which reflect specific aspects of user satisfaction, can enhance both the performance and interpretability of Review-based recommender systems. Tripadvisor  [^11] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/, for instance, collects ratings on different criteria including location, Service, Cleanliness, Value, Sleep Quality, etc. The ratings on the individual criteria represent users’ fine-grained preferences because not all the criteria are equally important to a user. Therefore, integrating these criteria rating information into the review and overall rating features could potentially improve the performance and provide better interpretability. Furthermore, multi-criteria rating information is valuable in mitigating the issue of data sparsity. In cases where reviews are sparse, leveraging multiple criteria can aid in learning user preferences. Recently, MRRRec (Hasan et al., 2022), and (Hasan and Ding, 2023) demonstrated the effectiveness of integrating multicriteria rating data with review to enhance the performance of the recommender system. However, a sophisticated approach is required to take full advantage of the utilizing criteri-rating information into reviews.
9.4. Addressing Data Sparsity
Despite the richness of review data, the issues of data sparsity (Grčar et al., 2006) and the cold start problem persist, where new users or items have insufficient interactions to generate accurate recommendations. Future research could explore novel ways to tackle these problems, possibly by leveraging user demographics, and item metadata, or by using transfer learning techniques where knowledge from one domain is applied to another. Additionally, unsupervised and semi-supervised learning (Learning, 2006) methods that can effectively utilize sparse data to find patterns could play a vital role in mitigating these challenges.
9.5. Ethical Considerations and Bias Mitigation
As recommender systems increasingly influence what information and products users are exposed to, ethical considerations become paramount. Future research must prioritize the development of mechanisms to identify and mitigate the biases in review-based recommender systems. This includes biases introduced by demographic factors, manipulation through fake reviews, or systemic biases inherent in the data collection process. Ensuring fairness, diversity, and representativeness in recommendations is essential for building equitable systems that serve a broad user base effectively.
9.6. Real-time Adaptation and Personalization
User preferences are not static; they evolve based on trends, seasons, and personal circumstances. Future review-based recommender systems should aim for real-time adaptation capabilities, where recommendations can dynamically adjust based on emerging trends in review data or changes in user behavior. This calls for the development of agile models that can update themselves frequently and efficiently, possibly through incremental learning or online learning approaches.
9.7. Cross-Domain Recommendations
Exploring ways to apply knowledge from one product area to another and creating a common system to understand user likes and item features are key steps to make recommendation systems better and more flexible across different types of products. Domain adaptation techniques help by using what the system has learned in one area to improve recommendations in another, finding similarities between them (Khan et al., 2017). At the same time, creating a single framework that captures what users like and item details across all areas can help make recommendations more consistent and relevant, no matter the product type. These methods could lead to better recommendations across various product categories, making suggestions more personalized and useful to users.
9.8. Large Language Models and Recommendations
Large Language Models (LLMs) can play a crucial role in generating recommendations by leveraging their natural language understanding capabilities to analyze customer reviews and generate insightful, relevant suggestions (Fan et al., 2023; Laskar et al., 2023a, b). Recent studies have explored the potential of LLMs in recommender systems, with a focus on prompting engineering (Xu et al., 2024; He et al., 2023). These studies have highlighted the effectiveness of LLMs in recommendation tasks, particularly in understanding context, learning user preferences, and generating relevant recommendations. In developing review-based recommendations, LLMs can facilitate sentiment analysis, key theme extraction, personalized recommendations, predictive analysis, and review summarization through methods like zero-shot learning (Kojima et al., 2022), demonstrations (Dang et al., 2022), or fine-tuning (Raza and Ding, 2023). This strategy capitalizes on the advanced natural language understanding capabilities of LLMs, offering customers insightful, targeted recommendations that enhance decision-making and satisfaction (Liu et al., 2023).
9.9. Responsible Recommendations
Most research on recommender systems has traditionally aimed at making the recommended results more accurate. However, the importance of ethical considerations and the impact of these systems on society has increasingly come into focus (Wang et al., 2022). There is a need for focusing on the aspects of privacy, diversity, and fairness, and how they interconnect with the notion of responsibility in aspect-based recommender systems (Elahi et al., 2022). While, the current literature often lacks a systematic exploration of these integrated approaches, particularly in the context of review-based recommender systems, these systems could greatly benefit from such responsible practices.
10. CONCLUSIONS AND FUTURE WORK
This paper presents a comprehensive exploration of review-based recommender systems, demonstrating their significant potential. We introduce various classification schemes based on feature extraction and review integration strategies. Through a detailed analysis of state-of-the-art methodologies, we show how leveraging user-generated reviews can substantially enhance the personalization and accuracy of recommendations, moving beyond the limitations of traditional rating-based systems. We identify the key challenges, including the complexity of learning user/item representation from review, review-integration strategies, the need for scalable solutions to handle the vast datasets, and the importance of ethical considerations in data handling. Despite these challenges, advancements in machine learning, deep learning, and NLP techniques offer promising avenues for extracting valuable insights from reviews, thereby enriching recommender systems with detailed understandings of user preferences.
Looking forward, the integration of multimodal data presents an exciting opportunity for further enriching review-based recommender systems, enabling a holistic approach to understanding user sentiments and preferences. Addressing data sparsity and the cold start problem remains a critical area for future research. Moreover, as recommender systems increasingly influence user choices and experiences, ethical considerations around bias mitigation and data privacy must be central to future developments. Ensuring that recommender systems are fair, and transparent, and respect user privacy will be paramount in fostering trust and acceptance among users.
11. Acknowlegements
This work is supported by the Natural Sciences and Engineering Research Council of Canada and the York Research Chairs program. Jimmy Huang (jhuang@yorku.ca) and Chen Ding (cding@torontomu.ca) are the contact authors of this paper. We would like to extend our gratitude to Syed Faizan for his invaluable assistance in gathering and organizing several papers